OpenStudy (anonymous):

x1=16, x2=0, y1=16 and y2=18 y2-y1/x2-x1=

OpenStudy (anonymous):

-1/8 = (18-16)/(0-16) Slope!!

OpenStudy (anonymous):

slope=-1/8